# A benchmark for computational analysis of animal behavior, using animal-borne tags

### v2

@@ Abstract
Animal-borne sensors (‘bio-loggers’) can record a suite of kinematic and environmental data, which can elucidate animal ecophysiology and improve conservation efforts. Machine learning techniques are used for interpreting the large amounts of data recorded by bio-loggers, but there exists no common framework for comparing the different machine learning techniques in this domain. To address this, we present the Bio-logger Ethogram Benchmark (BEBE), a collection of datasets with behavioral annotations, as well as a modeling task and evaluation metrics. BEBE is to date the largest, most taxonomically diverse, publicly available benchmark of this type, and includes 1654 hours of data collected from 149 individuals across nine taxa. In addition, using BEBE, we test a novel self-supervised learning approach to identifying animal behaviors based on bio-logger data, using a deep neural network pre-trained with self-supervision on data collected from human wrist-worn accelerometers. We show that this approach out-performs common alternatives, especially in a setting with a low amount of training data. Datasets, models, and evaluation code are made publicly available at https://github.com/earthspecies/BEBE, to enable community use of BEBE as a point of comparison in methods development.
[…]
@@
Animal behavior is of central interest in ecology and evolution, because an individual’s behavior affects its reproductive opportunities and probability of survival [17]. Additionally, understanding animal behavior can be key to identifying conservation problems and planning successful management interventions [4], for example in rearing captive animals prior to reintroduction [83], designing protected areas [76], and reducing dispersal of introduced species [78].
One increasingly utilized approach for monitoring animal behavior is remote recording by animal-borne tags, or bio-loggers [67, 86, 87]. These tags can be composed of multiple sensors such as accelerometer, gyroscope, altimeter, pressure, GPS receiver, microphone, and camera, which record time-series data on an individual’s behavior and their in situ environment. Additionally, bio-logger datasets can include data from many-hour tag deployments on multiple individuals.
@@
For classifying the behaviors underlying bio-logger data, researchers are increasingly using supervised machine learning (ML) techniques [81]. In a typical workflow, a human annotates some of the recorded bio-logger data with the tagged individuals’ behavioral states using a pre-determined ethogram, based on observations made simultaneously with data recording. These annotated data are used to train a ML model, which is then used to predict behavioral labels for the remaining un-annotated portion of the dataset. A test dataset, which is held out from the training stage, can be used to evaluate how well the trained model is able to perform this behavior classification task. In this manner, ML can help scientists to minimize manual effort required to ascribe behavioral labels to bio-logger data, or extend behavioral labels to data where manual ground-truthing is not possible.
@@
For behavior classification from bio-loggers, a benchmark could assess model performance across a breadth of study systems in order to identify relevant patterns, such as how modeling decisions can influence classification performance. Indeed, researchers have analyzed multi-species datasets to identify best practices for other key challenges in bio-logging, such as sensor calibration [28] and signal processing [46, 62]. To our knowledge, five previous studies have applied one or more behavior classification techniques on multiple bio-logger datasets, and the datasets used in these studies could hypothetically be used as a benchmark [11, 6, 88, 71, 85]. However, four of these do not make all of their bio-logger data publically available[11, 6, 88, 71] and therefore are not available for community use as a benchmark. The fifth does not reflect deployment variability in bio-logger data due to the inclusion of a single individual for each of four species [85].
@@
As a first application of BEBE, we make and test several hypotheses about ML usage in bio-logger data (Table 1). We base these hypotheses on recent trends in ML, as well as based on their potential to influence the workflow of researchers using ML with bio-logger data. First, in many applications of ML, deep neural networks that make predictions based on raw data, out-perform classical ML methods, which make predictions based on hand-crafted summary statistics or features [44]. In practice, the most commonly used ML methods for bio-logger data have been classical techniques such as random forests [77, Table 2], for which feature engineering is a key challenging step. Contrary to this practice, we predict the dominance of deep neural networks will carry over to bio-logger data (H1).
Second, we examine how data recorded for one species can be used to inform behavior predictions for a different species. Some previous works (e.g. [55, 57, 11]) have adopted a cross-species transfer learning strategy, by training a supervised model on one species and then applying this trained model to another related species. Our approach is different in that we use self-supervised learning. In self-supervised learning, a ML model (typically a deep neural network) is pre-trained to perform an auxiliary task on an unlabeled dataset. Importantly, training the model to perform this auxiliary task does not require any human-generated annotations of the data. In this way, self-supervised pre-training can make use of a large amount of un-annotated data that is easy to obtain (e.g., a species where a large dataset exists). Later, the pre-trained model can be trained (or fine-tuned) to perform the task of interest (such as behavior classification in a different species), using a small amount of annotated data. By learning to perform the auxiliary task the model learns a set of features, which often provide a good set of initial model parameters for performing the task introduced in the fine-tuning step. Inspired by the recent success of self-supervised learning in other domains of ML (e.g. language [10] and computer vision [12]), we predict that a deep neural network pre-trained on a large amount of human accelerometer data will outperform alternative methods, after fine-tuning (H2). Moreover, the self-supervised pre-training step can reduce the amount of annotated data required to meet a given level of performance [14]. Therefore, we predict that this trend will hold when we reduce the amount of training data by a factor of four (reduced data setting; H3).
@@
While we focus on a specific classification task in this study, all datasets, models, and evaluation code presented in BEBE are available at https://github.com/earthspecies/BEBE for general community use. Researchers may use the standardized task to test classification methods, or adapt BEBE datasets for their own research questions. Given that one aim of BEBE is to improve our understanding of classification methods in bio-logger data, we are also seeking contributions to create an expanded benchmark with improved taxonomic coverage, a broader range of sensor types, additional standardization, and a wider variety of modeling tasks. Details about how to contribute in this way can also be found at our GitHub.
@@
(H1) Deep neural network-based approaches will outperform classical approaches based on hand-chosen summary statistics.
Yes
(H2) Self-supervised pre-training using human accelerometer data will improve classification performance.
Partly
(H3) Self-supervised pre-training using human accelerometer data will improve classification performance when the amount of training data is reduced by a factor of four by removing individuals.
Yes
(H4) In terms of a single model’s predictive performance, there is minimal improvement in some behavior classes when increasing the amount of training data by four times by adding individuals.
Yes
@@ 1.1 Benchmark Datasets
We brought together nine animal motion datasets into a benchmark collection called the Bio-logger Ethogram Benchmark (BEBE) (Table 1.1). BEBE introduces a novel, previously unpublished dataset (Crow); otherwise, these data were all collected in previous studies. Of the datasets included in BEBE, four are publicly available for the first time (Whale, Crow, Rattlesnake, Gull) and five were already publicly available (HAR, Polar bear, Sea turtle, Seals, Dog).
@@
Table 2: Summary of datasets in BEBE. Out of nine datasets, one comes from humans, three come from other terrestrial species, three come from aquatic species, and two come from flying species. Datasets marked with an asterisk are publicly available for the first time in BEBE. Duration: the total duration of the dataset. Duration annotated: the duration of annotated data. Mean annot. dur.: the average duration an individual spends in a behavioral state (described in Methods). Annot. method: the type of data that was used to make the behavioral annotations. $C$: Creative Commons.
HAR
[2, 65]
[…]
[20]
Polar bear
[56, 57]
Dog
[40, 82]
[…]
[27]
Sea Turtle
[34]
Seals
[43, 42]
[…]
[39]
[…]
@@
Species
Human Western diamondback rattlesnake
Polar bear
Domestic dog Humpback whale
Green turtle
Otariid spp. Black-tailed gull Carrion crow
@@
100-300
.4-.6 .4-.6
Location
Genoa,
Italy
Texas, USA Beaufort Sea,
Arctic Ocean
Helsinki,
Finland
Wilhelmina Bay,
@@
Peninsula
Grande Anse
@@
France
West coast of
Australia
Kabushima Island,
Japan
León,
@@
position
Waist Body
Neck
Back and neck Dorsal surface or flank
Carapace
Back Back or abdomen Tail
# indiv.
30 13
5
45 8
14
12 11 11
@@
Walk
Moving
Not moving
Walk
@@
Gallop
Feed dive
@@
Rest dive
Swim
@@
channels
TIA
@@
TIA TIA
Conductivity
2x TIA
2x gyroscope
TIA
Depth
Speed
TIA
Gyroscope
Depth
TIA
Depth
TIA TIA
Duration (hrs)
6.2 30.9
1108.4
29.5 184.6
77.1
14.0 85.7 114.6
Duration annotated (hrs)
4.2 30.9
196.1
16.9 114.1
67.8
11.6 85.0 3.4
@@
method
Direct obs. Direct obs.
On tag
video
@@
video
On tag
video
On tag
@@
License
Custom $C$
Public
Domain
$C$ $C$
Public
Domain
$C$ $C$
$C$
In addition to the time series bio-logger data, each dataset in BEBE comes with human-generated behavioral annotations (Figure 2, colored bars). In each dataset, each sampled time step is annotated with the current behavioral state of the tagged individual, which can be one of several discrete behavioral classes. At some time steps, it was not possible to observe the individual, or it was not possible to classify the individual’s behavior using the predefined behavioral classes. In these cases, this time step is annotated as Unknown. These Unknown behavioral annotations are disregarded during model training and evaluation (also see Methods).
There are multiple time scales of behavior represented across the nine ethograms in BEBE, with some datasets including brief activities (e.g. shaking), and some including longer duration activities (e.g. foraging). In Table 1.1 we report the mean duration (in seconds) of an annotation in each dataset, as a rough estimate of the mean duration an individual spends in a given behavioral state.
In addition to Figure 2, we provide additional data visualizations in the Supplementary Information. Examples of each behavior class with the full set of channels reveal varying degrees of stereotypy in the behavioral classes (Supplementary Figures S1-S10). For example, in the Sea Turtle dataset, Stay at surface appears more stereotyped than Feed. Summary statistics across different individuals for each behavior class suggest the presence of discriminative features for some datasets, as well as differences between individuals (Supplementary Figures S11-S19). For example, in the Rattlesnake dataset, Move shows higher variance in raw accelerometer values compared to Not Moving, although to different degrees in different individuals.

### v3

@@ Abstract
Background: Animal-borne sensors (‘bio-loggers’) can record a suite of kinematic and environmental data, which are used to elucidate animal ecophysiology and improve conservation efforts. Machine learning techniques are used for interpreting the large amounts of data recorded by bio-loggers, but there exists no common framework for comparing the different machine learning techniques in this domain. This makes it difficult to, for example, identify patterns in what works well for machine learning-based analysis of bio-logger data. It also makes it difficult to evaluate the effectiveness of novel methods developed by the machine learning community.
Methods: To address this, we present the Bio-logger Ethogram Benchmark (BEBE), a collection of datasets with behavioral annotations, as well as a modeling task and evaluation metrics. BEBE is to date the largest, most taxonomically diverse, publicly available benchmark of this type, and includes 1654 hours of data collected from 149 individuals across nine taxa. Using BEBE, we compare the performance of deep and classical machine learning methods for identifying animal behaviors based on bio-logger data. As an example usage of BEBE, we test an approach based on self-supervised learning. To apply this approach to animal behavior classification, we adapt a deep neural network pre-trained with 700,000 hours of data collected from human wrist-worn accelerometers.
Results: We find that deep neural networks out-perform the classical machine learning methods we tested across all nine datasets in BEBE. We additionally find that the approach based on self-supervised learning out-performs the alternatives we tested, especially in settings when there is a low amount of training data available.
Conclusions: In light of these results, we are able to make concrete suggestions for designing studies that rely on machine learning to infer behavior from bio-logger data. Therefore, we expect that BEBE will be useful for making similar suggestions in the future, as additional hypotheses about machine learning techniques are tested. Datasets, models, and evaluation code are made publicly available at https://github.com/earthspecies/BEBE, to enable community use of BEBE.
Equal contribution
@@
1 Background
Animal behavior is of central interest in ecology and evolution because an individual’s behavior affects its reproductive opportunities and probability of survival [1]. Additionally, understanding animal behavior can be key to identifying conservation problems and planning successful management interventions [2], for example in rearing captive animals prior to reintroduction [3], designing protected areas [4], and reducing dispersal of introduced species [5].
One increasingly utilized approach for monitoring animal behavior is remote recording by animal-borne tags, or bio-loggers [6, 7, 8]. These tags can be composed of multiple sensors such as an accelerometer, gyroscope, altimeter, pressure, GPS receiver, microphone, and/or camera, which record time-series data on an individual’s behavior and their in situ environment. Additionally, bio-logger datasets can include data from many-hour tag deployments on multiple individuals.
@@ 1 Background
For classifying the behaviors underlying bio-logger data, researchers are increasingly using supervised machine learning (ML) techniques [11]. In a typical workflow, a human annotates some of the recorded bio-logger data with the tagged individuals’ behavioral states using a pre-determined ethogram, based on observations made simultaneously with data recording. These annotated data are used to train a ML model, which is then used to predict behavioral labels for the remaining un-annotated portion of the dataset. A test dataset, which is held out from the training stage, can be used to evaluate how well the trained model is able to perform this behavior classification task. Using the predicted behavioral labels allows large datasets to be leveraged to address scientific questions, for example, through estimating activity budgets that vary by time or individual [12] or by environmental conditions [13]. In this manner, ML can help scientists to minimize manual effort required to ascribe behavioral labels to bio-logger data, or extend behavioral labels to data where manual ground-truthing is not possible.
@@
For behavior classification from bio-loggers, a benchmark could assess model performance across a breadth of study systems in order to identify relevant patterns, such as how modeling decisions can influence classification performance. Indeed, researchers have analyzed multi-species datasets to identify best practices for other key challenges in bio-logging, such as sensor calibration [46] and signal processing [47, 48]. Previous studies have applied one or more behavior classification techniques on multiple bio-logger datasets, with varying degrees of variability in the species and individuals included [49, 50, 23, 51, 52, 53, 20, 54, 55], but none have attempted to compile a managed, publicly available and diverse database that others could compare against as a benchmark.
@@
As a first application of BEBE, we make and test several hypotheses about ML usage in bio-logger data (Table 1). We base these hypotheses on recent trends in ML, as well as based on their potential to influence the workflow of researchers using ML with bio-logger data. First, in many applications of ML, deep neural networks that make predictions based on raw data, out-perform classical ML methods such as random forests, which make predictions based on hand-crafted summary statistics or features [56]. Deep neural networks that operate on raw data, such as convolutional and recurrent neural networks, have previously been applied to behavior classification in wild non-human animals [49, 57, 58, 59], captive non-human animals [60] and in human activity recognition [61, 62, 63, 64, 65, 66, 67]. We distinguish these from the multilayer perceptron, a type of neural network that utilizes hand-crafted features and has been used in several bio-logging studies (e.g.,  [68, 14, 52]). Studies differ on whether deep neural networks show performance benefits compared to classical methods (e.g., [57, 49]), and random forests remain the most commonly used ML methods for bio-logger data [69, Table 2], for which feature engineering is a key challenging step. In line with ML trends and results from [49], we predict that deep neural networks will outperform techniques using hand-chosen features (H1).
Second, we examine how data recorded for one species can be used to inform behavior predictions for a different species. Some previous works (e.g. [49, 22, 23]) have adopted a cross-species transfer learning strategy, by training a supervised model on one species and then applying this trained model to another related species. Our approach is different in that we use self-supervised learning. In self-supervised learning, a ML model (typically a deep neural network) is pre-trained to perform an auxiliary task on an unlabeled dataset. Importantly, training the model to perform this auxiliary task does not require any human-generated annotations of the data. In this way, self-supervised pre-training can make use of a large amount of un-annotated data that is easy to obtain (e.g., a species where a large dataset exists). Later, the pre-trained model can be trained (or fine-tuned) to perform the task of interest (such as behavior classification in a different species), using a small amount of annotated data. By learning to perform the auxiliary task the model learns a set of features, which often provide a good set of initial model parameters for performing the task introduced in the fine-tuning step. Inspired by the recent success of self-supervised learning in other domains of ML (e.g. language [70] and computer vision [71]), we predict that a deep neural network pre-trained on a large amount of human accelerometer data [61] will outperform alternative methods, after fine-tuning (H2). Moreover, the self-supervised pre-training step can reduce the amount of annotated data required to meet a given level of performance [72]. Therefore, we predict that this trend will hold when we reduce the amount of training data by a factor of four (reduced data setting; H3).
@@
While we focus on a specific classification task in this study, all datasets, models, and evaluation code presented in BEBE are available at https://github.com/earthspecies/BEBE for general community use. Researchers may use the standardized task to test classification methods, or adapt BEBE datasets for their own research questions (see Discussion for examples). Given that one aim of BEBE is to improve our understanding of classification methods in bio-logger data, we are also seeking contributions to create an expanded benchmark with improved taxonomic coverage, a broader range of sensor types, additional standardization, and a wider variety of modeling tasks. Details about how to contribute in this way can also be found at our GitHub.
In summary, the main contributions of this study include:
  1. 1.
    Publicly available multi-species bio-logger benchmark dataset, centered on tri-axial accelerometers
  2. 2.
    Standardized evaluation framework for supervised behavior classification, with accompanying code and model examples
  3. 3.
    Demonstration of benchmark usage to investigate patterns in ML behavior classification performance (Table 1), including:
    • A comparison of deep learning and classical techniques on non-human bio-logger data
    • Successful cross-species application of a self-supervised neural network trained on human bio-logger data (based on [61])
@@
(H1) Deep neural network-based approaches will outperform classical approaches based on hand-chosen summary statistics. Yes, for the approaches we tested
(H2) Self-supervised pre-training using human accelerometer data will improve classification performance. Partly
(H3) Self-supervised pre-training using human accelerometer data will improve classification performance when the amount of training data is reduced by a factor of four by removing individuals. Yes
(H4) In terms of a single model’s predictive performance, there is minimal improvement in some behavior classes when increasing the amount of training data by four times by adding individuals. Yes
@@ 2.1 Benchmark Datasets
We brought together nine animal motion datasets into a benchmark collection called the Bio-logger Ethogram Benchmark (BEBE) (Table 2). BEBE introduces a previously unpublished dataset (Crow); otherwise, these data were all collected in previous studies. Of the datasets included in BEBE, four are publicly available for the first time (Whale, Crow, Rattlesnake, Gull) and five were already publicly available (HAR, Polar bear, Sea turtle, Seals, Dog). We summarize datasets’ hardware, data collection, ethogram definition, and ground-truthing in Section 2.1.2. For full details, including details on synchronization, calibration, and annotation validation, we refer readers to the original papers.
@@
{tblr}
width = colspec = Q[108]Q[81]Q[135]Q[87]Q[83]Q[100]Q[75]Q[108]Q[83]Q[67], roweven = Silver, hline1-2,18 = -, & Human Activity Recognition (HAR)
[41, 74] Rattlesnake
[17] Polar Bear
[40, 22] Dog
[29, 36] Whale
[38] Sea Turtle
[18] Seals
[39, 21] Gull
[35] Crow
@@
Species Human Western
diamondback
rattlesnake Polar bear Domestic dog Humpback whale Green turtle Otariid spp. Black-tailed gull Carrion crow
@@
100-300 .4-.6 .4-.6
Location Genoa,
Italy Texas, USA Arctic Ocean Helsinki,
Finland Wilhelmina Bay,
@@
Peninsula Grande Anse
@@
France Marine
facilities
on west coast
of Australia Kabushima Island,
Japan León,
@@
position Waist Body Neck Back and neck Dorsal surface or flank Carapace Back Back or abdomen Tail
# indiv. 30 13 5 45 8 14 12 11 11
@@
Walk Moving
Not moving Walk
@@
Gallop Feed dive
@@
Rest dive Swim
@@
channels TIA
Gyroscope TIA TIA
Conductivity 2x TIA
2x gyroscope TIA
Depth
Speed TIA
@@
Depth TIA
Depth TIA TIA
Hardware Galaxy S II
Samsung AXY-3,
AXY-4
(Technosmart) Video collar: Exeye
Logger: TDR10-X-340D
(Wildlife Computers) ActiGraph
GT9X Link
(ActiGraph LLC) DTAG [75] CATS
(Customized Animal
Tracking Solutions) CEFAS G6a+
(CEFAS Technology
Ltd) TDK MPU-9250 (InvenSense) miniDTAG [37],
TIA: KX022-1020 (Kionix)
Attach.
Method Belt Implant Collar Back: Harness
Neck: tape Suction Suction Fur seal: tape
Sea lion:
harness Tape, harness Elastic
Duration (hrs) 6.2 30.9 1108.4 29.5 184.6 77.1 14.0 85.7 114.6
Duration annotated (hrs) 4.2 30.9 196.1 16.9 114.1 67.8 11.6 85.0 3.4
Unknown behaviors (%) 33.4 0.0 82.3 42.7 38.2 12.1 17.0 0.8 95.7
@@
method Direct obs.,
@@
video On tag
video Direct obs.,
Off tag
video Motion,
on tag audio On tag
video Direct obs.,
Off tag
video On tag
video On tag
@@
License Custom Creative
Commons Public
Domain Creative
Commons Creative
Commons Public
Domain Creative
Commons Creative
Commons Creative
Commons